OpenStudy (anonymous):

-8(x-5)-2(5-x)=4x-80

OpenStudy (anonymous):

take out 2 from all parts..... -4(x-5)-(5-x)=2x-40

Parth (parthkohli):

Distribute. \(\Large \color{Black}{\Rightarrow -8x + 40 - 10 + 2x = 4x - 80 }\) \(\Large \color{Black}{\Rightarrow -6x + 30 = 4x - 80 }\) \(\Large \color{Black}{\Rightarrow -6x + 110 = 4x }\) \(\Large \color{Black}{\Rightarrow 110 = 10x}\) \(\Large \color{Black}{\Rightarrow x = 11 }\)

OpenStudy (anonymous):

so then -4x+20-5+x=2x-40 -3x+15=2x-40 15=5x-40 55=5x x=11